[03] Letan = n3x, then lim lan+1/an] = (A)x/31; (B) 3xl; (C)|3/x]; (D) xl; (E) None of these. [04] The radius R of convergePLEASE ANSWER AS MANY AS YOU CAN I NEED THE EXPLANATIONS TO STUDY ((((((
Answer #1

3) Ean=5nnxn we have an=namach Ponti (19+1) 39+cnti =(1+1)3 31032 Then lim I antil = lim an nyo 1 (n+1337.3 (869.a nga But- E o = (1+oja = 3 ROBO REŠ Hence Rading of convergent (R)= 1/2 ou The Interval of convergence of Enghoch :- o we have cx-ojn6 Criven power series Exe (n3) som { to t we have j anti = intus g9+1 I have an= ulan an= 2.3n. Then by Ratio Test we know= lot more J 14 m 1 - Nitoas Itin s1. . Hence Rading of convergoce (R321. Wie +20+2 > 5[1212)} *2012 ]: E[ mm] X2, we have I0 539/! Then by Ratio Test lims | ant la limo lenta lantil ein 8. 3 x n in l ndoo | lim hto huco nto 1 an lim. 1 =3 > Then by